### v1

# Dredge Word, Social Media, and Webgraph Networks for Unreliable Website Classification and Identification

Abstract
In an attempt to mimic the complex paths through which unreliable content spreads between search engines and social media, we explore the impact of incorporating both webgraph and large-scale social media contexts into website credibility classification and discovery systems. We further explore the usage of what we define as dredge words on social media—terms or phrases for which unreliable domains rank highly. Through comprehensive graph neural network ablations, we demonstrate that curriculum-based heterogeneous graph models that leverage context from both webgraphs and social media data outperform homogeneous and single-mode approaches. We further demonstrate that the incorporation of dredge words into our model strongly associates unreliable websites with social media and online commerce platforms. Finally, we show our heterogeneous model greatly outperforms competing systems in the top-k identification of unlabeled unreliable websites. We demonstrate the strong unreliability signals present in the diverse paths that users follow to uncover unreliable content, and we release a novel dataset of dredge words.
Dredge Word, Social Media, and Webgraph Networks for Unreliable Website Classification and Identification
Evan M. Williams  and Peter Carragher  and Kathleen M. Carley Carnegie Mellon University 5000 Forbes Ave. USA, Pittsburgh, PA, 15213-3890 {emwillia, pcarragh, carley}@andrew.cmu.edu
1 Introduction
On February 24, 2022, the day Russia began its invasion of Ukraine, Jacob Creech, a fringe Qanon conspiracy theorist, posted a series of unsubstantiated claims on Twitter. These tweets insinuated that the US had created COVID-19, and implied Russia’s invasion of Ukraine was actually being carried out to shut down US-funded “biolabs” in Ukraine and to prevent another global pandemic League (2022). Hours after the tweet, the conspiratorial website InfoWars published an article promoting the tweet, crediting its author for uncovering an “ulterior motive theory” League (2022). The tweet began circulating to other platforms as Google queries for “US biolabs” and “Ukraine Biolabs” spiked in the following days  [^1] https://trends.google.com/trends/explore?date=today%205-y&geo=US&q=Ukraine%20biolabs&hl=en. As only conspiratorial sites covered the tweet in its early hours, those would have likely been the only relevant search results until Snopes debunked the claim later that day Evon (2022).
In some ways, this is a success story of the current paradigm of combatting misinformation; fact-checkers responded rapidly to a quickly-spreading falsehood. In other ways, this highlights the challenges present in reactive misinformation interventions. Despite quick debunking by fact-checkers, the conspiracy was nonetheless echoed by established news sources like Fox News, and later amplified on the floor of the US SenateLeague (2022). Researchers later found that a coordinated campaign of social media accounts boosted the narrative on Twitter Alieva et al. (2022). As of 2024, despite the underlying claims having been debunked by many fact-checkers Chappell and Yousef (2022); Parachini (2022), a search engine audit study found that Google, Bing, and Yandex all returned URLs on their first page of results that promote the “Ukraine Biolabs” conspiracy Kuznetsova et al. (2024). This story also highlights an important and understudied dimension of misinformation spread: the interaction between social media and search engines.
An alternative to reactive fact-checking approaches are proactive algorithmic content moderation approaches. These can include modifying the recommendation and ranking systems to decrease the reach and virality of unreliable information sources. In search engines, this can mean downranking articles from unreliable domains, and on social media, this can mean ranking posts containing unreliable domains lower in user newsfeeds. For proactive approaches to succeed, platforms need systems that can rapidly detect and discover unreliable domains. Both classification and detection are important as unreliable websites can, and do, deploy a number of tactics to avoid blacklists Carragher et al. (2024). In this work, we present a novel approach to the tasks of unreliable domain detection and discovery that leverages signals from both large-scale social media data and webgraph data.
Directing users to search engines can be an effective tactic to spread misinformation as, since 2016, global surveys have consistently found that individuals report higher trust in search engines than traditional media Barometer (2024); McDuling (2015). Thus by pointing users to search engines, either explicitly or implicitly, bad actors can create a false sense of content reliability. Little is understood about these pathways, and research has only recently begun exploring the ways social media users end up in data voids. The concepts of problematic queries Golebiewski and Boyd (2019) and keyword signaling Tripodi (2019) are highly related, but have only been explored in case studies. The recently-proposed concept of search directives provides a clear and observable path that could direct users to unreliable content on search engines. A search directive is content that is explicitly intended to prompt an online search, e.g., user $a$ directs user $b$ to "look up Chemtrails on Google" Robertson et al. (2023). In the context of data voids, search directives as a concept have two limitations: First, while search directives can direct users to unreliable content, they can also direct users to innocuous or helpful information. Second, as the search trends for “Ukraine Biolabs” demonstrate, users can be driven to unreliable content through search engines by querying phrases that appear without explicit search directives. To help fill these gaps, we propose the concept of dredge words—terms or keyphrases for which unreliable domains rank highly on search engines.
By attempting to explicitly incorporate each of the paths that users take to unreliable websites into GNNs, we seek to explore the dynamics that connect the spread of unreliable content on social media and search engines. First, we attempt to demonstrate domain (un)reliability signal present in large-scale social media data and small-scale webgraph data with varying levels of context. With respect to website credibility classification, we use graph neural networks to demonstrate the domain reliability signal in user-domain interaction networks on large-scale twitter data, on small-scale webgraph networks, and homogeneous and heterogeneous combinations of the two. Through comprehensive graph neural network ablations, we demonstrate that heterogeneous graph models that leverage context from both webgraphs and social media data outperform homogeneous and single-mode approaches. We additionally demonstrate that designing a training curriculum based on relative website reliability scores further increases model performance.
Finally, we provide the first exploration of dredge words on social media—terms or keyphrases for which unreliable domains rank highly in search engines. We demonstrate the usefulness of these dredge words in an unreliable domain classification and discovery process that mimics how social media users may transition from social platforms to the consumption of misinformation via search engines. Surprisingly, we also find that dredge words frequently surface social media URLs in top SERP positions, thus providing a bidirectional path that often leads back to social media. We show our heterogeneous model greatly outperforms competing systems in the top-k discovery of unlabeled unreliable websites. We publicly release the code, webgraph data collected for this project, a dataset of 3,939 dredge words for 46 unreliable news domains, and the resulting SERPs for each of the dredge word queries [^2] Artefacts will be publicly released following review period.
2 Related Works
2.1 Social Media
Classifying individual texts or articles is a core problem addressed by research in misinformation detection. Such detection methods have typically relied on website content and social media data Castelo et al. (2019); Chen and Freire (2020); Silva et al. (2021). Castelo et al. proposed a topic-agnostic detection system Castelo et al. (2019) that identifies unreliable articles based on Linguistic Inquiry and Word Count features. Chen and Freire adopted the method for the task of unreliable domain discovery Chen and Freire (2020). Their discovery system capitalized on user tendencies within the social graph, wherein a user who tweets a URL from a known unreliable source is likely to also tweet URLs from yet unknown sources. Similarly, Silva et al. combine website content and social context resulting in a misinformation page detection method that leverages heterogeneous data types Silva et al. (2021), with a focus on early detection across a broad range of topics.
2.2 Webgraphs
The predictive power of webgraphs and social media data have been demonstrated separately in several contexts. Aswani et al. detect SEO manipulation by link-building sites by clustering on Pagerank score and Domain Authority Aswani et al. (2021). For detecting news site bias, Aires et al. Patricia Aires et al. (2019) scrape cross-links from a list of prominent news sites from Media Bias Fact Check (Zandt, 2022). Sehgal et al. explore a case where misinformation was spread through coordinated hyperlink and social media networks Sehgal et al. (2021) on a dataset of 1.4k URLs. Another recent case study demonstrated the manipulated webgraph linkages of unreliable pseudo-thinktanks Williams and Carley (2023). More recently, Carragher et al. (2024) proposed a webgraph-based model for unreliable domain detection and discovery tasks. For discovery, the authors use backlinking domains in the webgraph in a snowball sampling approach. The authors demonstrate that GNN models trained on SEO attributes and webgraph data proved effective for detection.
2.3 Combining Webgraphs and Social Media Context
Zhang and Cabage show that social media and SEO promotion have different strengths; they find that social-sharing results in immediate but short-term boosts to traffic, while the benefits of link-building are slower to realize, but last longer Zhang and Cabage (2017). This hints that a combination of both webgraph and social media data should be used in investigating misinformation sources. However, to our knowledge, this is the first work that attempts to combine the two contexts for unreliable domain classification or discovery.
3 Data
We construct datasets based on the paths that users can take to unreliable websites. In our social media and webgraph data, we attempt to capture explicit paths users take to unreliable websites—i.e. clicking a URL. The dredge words we collect for a subset of the unreliable domains represent implicit paths.
3.1 URL Labels
We use the domain credibility labels published by Lin et al. Lin et al. (2023). The authors of the work align the domain reliability ratings of all domains ranked by 6 expert groups and run imputation followed by principal component analysis to generate aggregate ratings for 11,520 domains Lin et al. (2023). To our knowledge, this is the most comprehensive public domain reliability rating list as of spring of 2024. The domains in this list are ranked by the calculated first principal component where low scores correspond to high expert agreement on unreliability and high scores correspond to expert agreement on reliability. While these scores are very sensible across wide principal component score gaps, locally the relative orderings are less clear. Treating website reliability as continuous also muddies the interpretability and the discovery process, as there’s not a transparent reason a website should have a score of 0.571 as opposed to 0.570. To address this issue, we elected to binarize the website reliability labels. After manual inspection of the data, we elected to consider the bottom two quintiles as unreliable (corresponding to a principal component threshold of 0.5162).
3.2 Webgraph and features
As extracting complete webgraph information for these domains is computationally challenging, researchers have generally needed to use third-party tools to extract local webgraph attributes. We use the SEO toolkit service Ahrefs333ahrefs.com to extract the 10 domains which link to each target domain at the highest volume (the highest-volume back-linking domains). We chose Ahrefs because researchers have previously shown the usefulness of its data in domain reliability classification and discovery applications Carragher et al. (2024); Williams and Carley (2023). We were unable to pull backlinks for 193 of the domains. A random sample of 10 of 193 domains and found 9 of them were either dead or unreachable. We therefore elected to drop these 193 domains, leaving us with 11,327 unique reliability-labeled domains and 32,431 unique unlabeled backlinking domains for a total of 43,758 domains. For each of these 43,758 domains, we pull 23 attributes, which for an individual domain, contains fields like the total number of backlinks and outlinks, number of backlinks coming from .edu or .gov domains, and number of referring pages  [^4] https://ahrefs.com/api/documentation/metrics.
Similar to the backlink network constructed in Carragher et al. (2024), there is clear assortativity across label-reliability groups as can be seen in Figure 1. The prevalence of unlabeled nodes in our network overwhelms a node attribute assortativity coefficient calculation $r_{A}=-0.336$. We re-calculate node attribute assortativity on an induced subgraph that only maintains edges between labeled nodes. The subgraph contains less structure and 176 components, but nevertheless displays strong positive associativity $r_{A}=0.376$. In other words, domains of the same (binary) reliability are more likely to link to one another than linking to a domain of a different reliability. This provides a strong justification for the use of network-based models.
Figure 1: Webgraph colored by domain reliability labels. The network contains 6,861 reliable (blue) websites, 4,466 (red) unreliable websites, and 32,431 unlabeled (grey) backlinking websites.
3.3 Twitter Data
For social media context, we use a Twitter dataset constructed by querying COVID-related keywords555coronavirus, Wuhan virus, Wuhanvirus, 2019nCoV, NCoV, NCoV2019, covid-19, covid19, covid 19 via Twitter’s streaming API between January 29, 2020 and June 26, 2022. Due to server issues and API limitations, 121 days over the time period have partial or missing data. However, these gaps are spread relatively evenly over the time period, and still provide strong coverage. Our final Twitter dataset contained 3.6 billion extracted tweets.
From this set of 3.6 billion tweets, we pull all Tweets that mention or link to one of our 11,520 websites, resulting in a dataset of 320M tweets that link to over 840k unique domains. Of the original 11,520 labeled domains, only 5,504 appeared within the Twitter data. To reduce the noise in the dataset we undertake several cleaning operations. Firstly, we drop all tweets that do not explicitly include a link to the URL. Second, users with fewer than 10 observed tweets over the time period were dropped. To further condense the dataset, we only include tweets which appeared at least 3 times in the dataset. As reposting and retweeting are important influence metrics on Twitter, this drops the tweets that likely did not receive as much attention. We also dropped domains that were only linked to by a single user and users that only linked to a single domain, as we hypothesized these pendulum nodes would be of limited use given the size of the graph. As a result of the expense of the webgraph attribute API, we chose to further restrict the Twitter data to only include tweets that mention at least one of the 43,758 domains for which we extracted attributes. This corresponds to keeping tweets that mention at least one of the 11,327 labeled domains for which we have attributes or tweets that mention one of the 11,327 and co-mention any of the 32,431 unlabeled domains. Following these cleaning operations, we are left with 555k users and 4.9M unique tweets which we observed tweeted or retweeted 91.2M times. This reduced dataset contains mentions of 2,475 reliability-labeled domains and 714 unlabeled back-linking domains.
3.4 Dredge Words
With API constraints in mind, for the 100 websites with the lowest PC1 score rank in Lin et al. (2023). We extract the top-ranked $k\leq 1,000$ Google keyphrases from Ahrefs, for a total of 34,646 keyphrases. We then used WebSearcher Robertson and Wilson (2020) to query each of these keyphrases on Google and extracted the first 10 URLs returned in each SERP. We kept all queries for which the target unreliable domain was returned in the top 10 Google search results. This resulted in our set of 3,939 dredge words spanning 46 unreliable domains. We created a separate query to find mentions of these keyphrases in the 3.6B tweet covid dataset. This yielded 5.7 million tweets containing dredge words. Many of the most common mentions were explicit mentions of domain names or organization names (which can be identical to twitter handles—like “gatewaypundit”, “infowars”, and “nvic” all received over 10,000 mentions. To filter down the data, we use regular expressions to ensure each drudge word begins with a hashtag, starts a tweet, or is preceded by a white space. This retains 213 dredge words in 421k tweets that qualitatively contain less noise. Of these 421k tweets, only 9,788 (2%) explicitly linked to the unreliable domain associated with the dredge word.
Figure 2: The top search results for the dredge word “silent assassination through amplified neurons”. The query surfaces fringe reddit subreddits followed by “beforeitsnews”, an unreliable news source.
Some examples of dredge words include “psychic attack” (28 Twitter mentions), “akashic record” (12), “flu shot injury” (6), and “fallcabal” (5). While the majority of the dredge words are in English, other languages are present in the list, including Chinese, Hindi, and Arabic. We use this condensed twitter dataset and the set of SERP resultss they yielded in our dredge-word-based unreliable domain discovery process. While we were initially interested in paths from social media to search engines, we find dredge word SERPs surprisingly demonstrate a strong paths to social media. Youtube was by far the most commonly-returned domain (4,304 times). Wikipedia, Reddit, Quora, Twitter, Amazon, and Facebook were also among the most commonly-returned domains, in part due to the widespread commercialization of pseudo-scientific concepts. For a more in-depth analysis of dredge words see Appendix A.
4 Methods
In order to evaluate the impact of additional levels of context on our models, we construct graphs that capture increasingly-granular levels of context present in the data.
4.1 Single Mode Graphs
We define two single-mode graphs to evaluate the effectiveness of the models using only webgraph data and only Twitter data. Let $\mathcal{W}$ be the set of all labeled and unlabeled websites $\{w_{1},w_{2},\dots,w_{n}\}$ contained in the extracted webgraph data and define $\mathcal{U}$ as the set of all Twitter users $\{u_{1},u_{2},\dots,u_{n}\}$ that link to domains $\in\mathcal{W}$. For our webgraph data, we define $\mathcal{G}_{w}=(\mathcal{W},A,Z,Y)$ where $A$ is a binary undirected adjacency matrix where a $1$ or 00 denotes the presence or absence of a connection between website $i$ and website $j$. $Z$ is a node feature matrix, and $Y$ is the label assigned to each node ($1$ = reliable, 00 = unreliable, $-1$ = unlabeled). We define a corresponding graph for Twitter user-to-website connections $\mathcal{G}_{t}=(\mathcal{U},A,Z,Y)$ where nodes are users and websites $\in\mathcal{U}$ and the set of edges captured by $A$ denote observed edges between $\mathcal{U}$ and $\mathcal{W}$.
In the $\mathcal{G}_{w}$, the constructed webgraph, we define $Z$ as the domain-level attributes extracted from Ahrefs. Following Carragher et al. (2024), we log-normalize the features of each domain. However, the social media network, $\mathcal{G}_{t}$ presents a challenge in defining $Z$, as users do not have website-level attributes. For this graph, we therefore define $Z_{n}$ as features calculated using using Node2Vec Grover and Leskovec (2016) parameterized with a walk length to 20, context size to 10, and 10 walks per node, and an embedding dimension of 23, chosen because it is identical to the number of features that we extract for each website from Ahrefs. $\mathcal{G}_{t}$ presents an additional challenge in that not all websites $\in\mathcal{W}$ are mentioned by users $\in\mathcal{U}$. We include social-media-only baselines, but these are trained and evaluated on only a network of user interactions with the 2,475 labeled domains mentioned in the social media data, which make these results not directly comparable with our other models.
4.2 Homogeneous Graph
We construct a homogeneous graph $\mathcal{G}_{Hom}=\mathcal{G}_{w}\cup\mathcal{G}_{t}$ where $\cup$ denotes a graph union operation. As node features for users and websites are different, we consider only features derived from the graph structure in the homogeneous model. Again, as features are different for users and websites, we calculate features using node2vec over the graph $\mathcal{G}_{Hom}$ using the same parameters as for the single-mode graph.
4.3 Heterogeneous Graph Construction
We construct a heterogeneous graph $\mathcal{G}_{Het}=(\mathcal{W}\cup\mathcal{U},A,Z,\phi,Y)$ where $\phi$ denotes a type of relation: websites-to-websites or users-to-websites. As a previous work has demonstrated the effectiveness of webgraph features in reliability classification Carragher et al. (2024), we consider only two different feature combinations. We consider a heterogeneous graph $\mathcal{G}_{Het}$ where websites have log-normed attributes from ahrefs and users are assigned the node2vec attributes calculated for $\mathcal{G}_{Hom}$. We consider a second heteroegeneous graph, $\mathcal{G}_{Het,u}$ where each user is assigned attributes extracted from their text content. For each user, we randomly sample, without replacement, 10 tweets for each observed user and embed them using multilingual distilBERT Sanh et al. (2019) [^6] All models in this paper were run or trained on a single NVIDIA GeForce RTX 3080.. This is a naive approach that excludes a massive amount of context for some users, particularly as we observed two users tweet over 1 million times in our filtered dataset. Clearly, only considering 10 tweets is suboptimal, but it is a standard practice in user embedding literature Pan and Ding (2019). A clear path for future work is to explore the impact more advanced user embedding strategies within this system.
4.4 Heterogeneous Dredge-word Graph Construction
We construct an additional heterogeneous graph $\mathcal{G}_{Het-d}$ that attempts to model the path through which dredge words can drive users to unreliable content on search engines. $\phi$ in this graph can denote connections between dredge words and websites extracted from SERPs or connections between dredge words and users $\in\mathcal{U}$. Dredge words embeddings were extracted using multilingual distilBERT, and these embeddings are used as features for dredge word nodes. Though the dredge word Twitter dataset was collected separately, but we choose to restrict the users we consider to those who also shared links to at least one known unreliable domain to facilitate model comparability. We finally define $\mathcal{G}_{Het-Dredge}$ as $\mathcal{G}_{Het-d}\cup\mathcal{G}_{Het}$.
4.5 Graph Neural Networks
A label-stratified 80/10/10 split on labeled websites is used to create training, validation, and test sets. For our one-mode, heterogeneous, and homogeneous experiments we use a simple 2-layer graph neural network using GraphSAGE convolutions proposed in Hamilton et al. (2017). We ran experiments using graph attention networks Veličković et al. (2017) and heterogeneous graph transformers Hu et al. (2020), but found marginal accuracy gains at the cost of longer training times. Consequently, we selected GraphSAGE layers. Each model consists of a SAGEConv layer with dropout, with a hidden dimension of 512, followed by a ReLU activation and a second SAGEConv layer with a log softmax activation function. For each model, we train for a maximum of 1,000 epochs with early stopping based on validation loss and a patience of 50. We use an Adam optimizer with a starting learning rate of 1e-3 and a cosine annealing learning rate scheduler. The heterogeneous GNN, the most complex network we consider, contains 3.2M parameters, and no model exceeded 5 minutes of training time to reach our early stopping convergence condition.
4.6 Curriculum Learning
We assume that highly reliable and highly unreliable websites are easier to differentiate in webgraphs than those that are mixed, mostly reliable, or mostly unreliable. A manual examination of the labeled URL data leads us to believe that this assumption is very reasonable when website content is considered (see appendix B).
We implement a slightly-modified version of the "Baby Steps" learning curriculum proposed and explored in Spitkovsky et al. (2010); Cirik et al. (2016). As domain labels from Lin et al. (2023) contain unified principal component scores of expert ratings of domain reliability, we can develop a curriculum that first learns labels of extremely reliable and extremely unreliable domains, and that gradually works from the extremes towards the reliability boundary. Using the original labeled domain dataset $\mathcal{D}$, we calculate quartiles of reliable labels and unreliable labels using principal component scores, and define these as an ordered set of batches $\{d_{1},d_{2},\dots,d_{10}\}$, where $d_{1}$ is the first quintile of the most reliable domains and $d_{10}$ is the fifth quintile of the unreliable domains. When using curriculum learning, we begin training the model $\mathcal{M}$ using $\{d_{1},d_{10}\}$, and following convergence, the model considers $\{d_{1},d_{10}\}\cup\{d_{2},d_{9}\}$. The model continues in this fashion until it converges on all available data. In our Baby Steps implementation, the we define model’s corresponding convergence as 10 epochs without an improvement in validation loss. We note that this is the curriculum convergence condition, and not the convergence condition of the model. More generally, for a website reliability curriculum $\mathcal{C}$, an even-length dataset $\mathcal{D}$, and a model $\mathcal{M}$, our implementation of the Baby Steps curriculum can be expressed as the procedure in Algorithm 1.
Algorithm 1 Modified Baby Steps Curriculum
$\mathcal{M}$, $\mathcal{D}$, $\mathcal{C}$
sort($\mathcal{D}$, $\mathcal{C}$)
$B\leftarrow\emptyset$
for $i\in 0,1,\dots\frac{k}{2}$ do
   while $\mathcal{M}$ not converged do
      $B\leftarrow B\cup\mathcal{D}[i]\cup\mathcal{D}[\frac{k}{2}-i]$
      $\mathcal{M}(B)$
return $\mathcal{M}$
4.7 Unreliable Domain Discovery
[…]
We compare our discovery processes with the webgraph-based discovery (WG-BD) process proposed in Carragher et al. (2024) and with the social media-based discovery process (SM-BD) proposed in Chen and Freire (2020). Specifically, we consider Precision@5, @10, @20, and we consider the partial F1 metric proposed in Chen and Freire (2020). This means we run the discovery process twice, once on the full domain list where we evaluate results manually with top-10 and top-20 accuracies, and again on a restricted domain list to compute the partial F1 metric. We additionally report partial precision and recall. We note that dredge words were only considered for 46 of the least reliable websites, so their inclusion in models is likely to harm the models, as those features only exist for a small set of the 11,327 labeled domains. We nonetheless report statistics and discovery evaluations.
Due to a lack of ground truth labels for evaluating newly discovered domains, Partial F1 measures the ability of the discovery system to find unreliable domains with respect to two known lists of unreliable sources, a seed list and an evaluation list. In this case, the seed list is the PoliticalNews dataset and the evaluation list is drawn from unreliable MBFC domains (Chen and Freire, 2020). As such, partial F1 measures the ability of a discovery system seeded on PoliticalNews to discover as many MBFC domains as possible (high partial recall), without discovering domains that are not in the MBFC list (high partial precision) (Castelo et al., 2019). We discuss the limitations of the partial F1 metric in more detail in Limitations and Appendix C.
5 Results
5.1 Credibility Classification
We present Accuracy and F1 statistics for GNN ablations in Table 1. $\mathcal{H}$ denotes a homogeneous GNN and $\mathcal{E}$ denotes a heterogeneous GNN. The subscripts $u$ and $w$ denote the use of the user network ($u$) and/or the webgraph network ($w$). We use graph-specific ($\mathcal{G}_{w}$, $\mathcal{G}_{t}$, or $\mathcal{G}_{w}\cup\mathcal{G}_{t}$) Node2Vec features in all GNNs without $f$ or $v$ subscripts. $f$ denotes the inclusion of webgraph features, $v$ denotes the inclusion of embedded user text, and $d$ denotes the model was trained on $\mathcal{G}_{Het-Dredge}$. Finally, $c$ denotes the inclusion learning curriculum in training. We note that the two social-media-only models, $\mathcal{H}_{u}$ and $\mathcal{H}_{u,c}$ are trained and evaluated on only the 2,475 labeled domains that were mentioned in the cleaned social media data.
We find that the heterogeneous model $\mathcal{E}_{u,w,c}$, which incorporates the user network, the webgraph, and curriculum learning outperforms all other models with an F1 of 0.7777±.003plus-or-minus0.7777.0030.$77$003. While the accuracy of the Social Media models is higher, those models are run only on a subset of the data, and thus are not directly comparable.
Model Accuracy F1
$\mathcal{H}_{w}$ 0.7686 ± .004 0.7545 ± .005
$\mathcal{H}_{w,f}$ 0.7799 ± .002 0.7714 ± .002
$\mathcal{H}_{w,f,c}$ 0.7791 ± .002 0.7705 ± .002
$\mathcal{H}_{u}^{*}$ 0.8012* ± .005 0.7443* ± .008
$\mathcal{H}_{u,c}^{*}$ 0.8113* ± .007 0.7582* ± .009
$\mathcal{H}_{u,w}$ 0.7523 ± .005 0.7654 ± .005
$\mathcal{H}_{u,w,c}$ 0.7696 ± .006 0.7559 ± .006
$\mathcal{E}_{u,w}$ 0.7797 ± .010 0.7707 ± .012
$\mathcal{E}_{u,w,c}$ 0.7865 ± .002 0.$7$003
$\mathcal{E}_{u,w,f,v}$ 0.7827 ± .007 0.7741 ± .008
$\mathcal{E}_{u,w,c,f,v}$ 0.7738 ± .118 0.7657 ± .018
$\mathcal{E}_{u,w,c,f,v,d}$ 0.7787 ± .004 0.7675 ± .007
Table 1: Mean Accuracy and F1 and standard deviations for each GNN ablation over 10 runs. $\mathcal{H}$ denotes homogeneous and $\mathcal{E}$ denotes a heterogeneous and subscripts denote included networks and features. * denotes social-media-only—these were run and evaluated on the 2,475 labeled domains mentioned in the Twitter data and thus are not directly comparable with other models.
5.2 GNN Discovery
Two annotators77725-30 year-old PhD candidates studying misinformation independently annotated the reliability of the top 20 predictions (sorted by prediction confidence) of the top performing model, $\mathcal{E}_{u,w,c}$, over the set of all unlabeled domains. Inter-annotator agreement in the first round had a Krippendorff’s $\alpha=0.78$. The annotators then met and resolved all disagreements. We compare our model with the webgraph-based discovery (WG-BD) approach in Carragher et al. (2024). As that paper’s discovery process did not rank discovered domains, we sort results by the misinformation classifier proposed in the work. Again, two independent annotators ranked the results, yielding a Krippendorff’s $\alpha=0.69$. For comparison with previous unreliable domain discovery approaches, we additionally report partial F1, defined in Chen and Freire (2020) as the set of true positive unreliable domains with credibility labeled “mixed” or worse by Media Bias Fact Check. Thresholding at a prediction confidence level of 0.7 (see Appendix C for sensitivity analysis), the Partial F1 of our top-performing GNN model is 0.25, which is largely in-line with Chen and Freire (2020) (0.29) and Carragher et al. (2024) (0.28). We observe that $\mathcal{E}_{u,w,c}$ outperforms both competing systems at each level of precision we consider.
5.3 Dredge Word Discovery
[…]
Model P@5 P@10 P@20 PF1 PP PR
$\mathcal{E}_{u,w,c}$ 1 1 0.9 0.25 0.18 0.35
$\mathcal{E}_{u,w,c,f,v,d}$ 0.07 0.04 0.26
SM-BD - - 0.7 0.29 0.24 0.37
WG-BD 0.2 0.5 0.65 0.27 0.24 0.31
[…]
6 Conclusion
We propose a model for unreliable domain detection that incorporates many of the diverse paths that users take to unreliable domains. We propose the idea of Dredge Words and highlight their bidirectional connections with social media. Finally, we demonstrate that our best-performing proposed model outperforms competing systems on the task of unreliable domain discovery.
Acknowledgements
The research for this paper was supported in part by the Office of Naval Research, MURI: Persuasion, Identity, & Morality in Social-Cyber Environments under grant N000142112749 and by the Knight Foundation. It was also supported by the center for Informed Democracy and Social-cybersecurity (IDeaS) and the center for Computational Analysis of Social and Organizational Systems (CASOS) at Carnegie Mellon University. The views and conclusions are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ONR or the US Government.
7 Limitations
While we attempted to mitigate limitations where possible, several substantial ones exist. First, the Twitter data were collected from 2020-2022, but the dredge words used in this paper were based on Google SERPs in 2024. We cannot say that the dredge words we extracted surfaced the same content in 2024 as when the phrases were initially used on Twitter. However, given the conspiratorial, niche, and pseudo-scientific nature of many of the dredge words we identified, we speculate that they would be unlikely to surface substantial reliable content, particularly as reliable sites have had several years to attempt to debunk—and thereby rank for—dredge word queries. Future work could explore the impacts of better temporal alignment, which would have the added benefit of covering data voids formed around breaking news stories.
The Partial F1 metric we use for discovery evaluation is limited, as previous work has shown many PoliticalNews and MBFC sites on which the metric relies are dead or no longer active Carragher et al. (2024). However, to our knowledge, no better metric exists for evaluating unreliable domain discovery systems. This is a needed avenue for future work. Additionally, the current process for extracting dredge words relies on paid third-party services. While we validated each of the extracted dredge words to ensure they are ranking in Google’s top-10, approaches to extract keyphrases from, for example, common-crawl webgraphs would provide more democratization of this line of research. We publish all dredge word data collected for this project in line with that goal.
Finally, dredge words were only extracted for 46 of the least reliable 11,327 labeled domains. Their inclusion in our models is therefore not particularly helpful. We elected to include the results from these models because they point to an interesting and understudied phenomenon: bidirectional paths between social media and search engines, which appear to be often targeted by monetary interests. In future work, we plan to repeat this process with dredge words extracted for a larger set of unreliable domains, along with keywords extracted from reliable domains.
8 Ethical Considerations
To avoid unintentionally boosting the search rankings of unreliable domains, we do not directly link to any referenced unreliable domains. For unreliable sites, we link to either archive.org or third-party reliability assessments of those domains. We further acknowledge the inherent challenges in determining the “reliability” of websites, and note that even reliable websites can occasionally publish inaccurate content. We defaulted entirely to third parties in our original acquisition of labels, but emphasize that any reliability-labeling ontology should be scrutinized to minimize harms. We would argue that the system we proposed could be generalized to any reasonable reliability ontology.
Acknowledging that the Twitter data likely contains examples of “sensitive events” which could compromise the privacy or safety of users, we adhere to Twitter’s terms of service and do not release this data [^8] https://developer.x.com/en/developer-terms/agreement-and-policy. Finally, we acknowledge that by publicly releasing these dredge-word lists, bad actors could use these terms to direct people to unreliable information. However, once data voids are identified, they are often plugged Norocel and Lewandowski (2023). We hope that by releasing this data, reliable websites can target these phrases to reduce both the visibility and the ease of access of unreliable content.
Appendix A Dredge Word EDA
A.1 Dredge Word SERPs
To understand the distributions of websites surfaced in SERPs, we will provide an illustrative example. One could imagine a user browsing a social media platform and coming across a tweet or post like the one in Figure 3.
Figure 3: A tweet about “indigo children”.
A user encountering new, potentially perplexing information, might then be curious and compelled to search “indigo children meaning” or just “indigo children” on Google—both phrases are dredge words. We provide the results for former query in Figure 4. We note that Google’s definition snippet expresses pseudo-scientific concepts. YouTube videos, followed by a Reddit post from the /r/Psychic subreddit were the fourth-to-the-seventh highest-ranked URLs when we manually searched this dredge word phrase in June of 2024. Another user who may have searched only “indigo children”, would have seen a snippet from Wikipedia which correctly calls indigo children pseudoscientific. However, the user would have seen four books for sale about indigo children alongside the snippet, and the second highest-ranking return was an article from an unreliable domain identifying the “13 signs you’re an indigo child”. Were the user to then scroll down to the 6th result, they would see an Amazon link to a book on indigo children.
Figure 4: SERP results for “Indigo Children Meaning”
A.2 Dredge EDA
We applied the python library langdetect Danilák (2014) to the list of 3,933 dredge words used in the paper, but we observe that automated language extraction of dredge words is a challenge. We observed many dredge words are disembodied fragments, names of people, or simply foreign words adopted by various communities. For example, several Hindi-language-origin yoga poses appear in the dredge word list: parivrtta anjaneyasana and parivrtta utkatasana. These terms surface gaia, a conspiracy pseudoscience website [^9] https://mediabiasfactcheck.com/gaia/. Some websites rank highly for the same terms in multiple langauges; the arabic and chinese words for (coral) both surface creation, a website that frequently promotes pseudoscience [^10] https://mediabiasfactcheck.com/christian-ministries-international/. Qualitatively, hatespeech is fairly uncommon in the dredge words we identified; 11 unique dredge word phrases contain the non-euphemistic writing of the n-word and 3 others contain a homophobic slur.
Appendix B Curriculum EDA
Among the websites rated most reliable, the style, content, and widespread name recognition would very likely result in most annotators to correctly identifying the sites as reliable (e.g., reuters.com, nasa.gov, smithsonianmag.com, nature.com). The landing pages of some of the least reliable labeled domains are also typically easy to categorize, with websites that contain toxic green backgrounds accompanied by blurry photos of George Soros  [^11] https://web.archive.org/web/20240112053008/http://www.endgamethemovie.com/, proclamations of white nationalism  [^12] https://web.archive.org/web/20240108044820/https://www.stormfront.org/forum/, or full links to alien siting websites and paid psychic services  [^13] https://web.archive.org/web/20240116140224/https://rense.com/. Though our webgraph model is content-agnostic, we hypothesize that the webgraph and site attributes will exhibit similarly "easy" patterns across the extremes. To test this hypothesis, we implement a curriculum learning batching procedure Bengio et al. (2009).
Appendix C Partial F1 threshhold sensitivity
To investigate the partial F1 metric we experiment with various classifier confidence thresholds. Naturally, as confidence increases, we see higher precision and lower recall (Figure 5). The disparity between partial recall and partial precision implies that our discovery process is not very precise. However, in the manual evaluation of Precision@k, we find precision is actually very high. This indicates a fundamental issue with the partial precision metric; it underestimates model performance, particularly as the MBFC reference list ages. This is in line with previous findings that unreliable domain lists quickly become outdated (Carragher et al., 2024), and it complicates the evaluation of discovery processes. Further work on domain reliability discovery metrics is needed.
For Precision@k assessments, two annotators were asked “is this website unreliable” and asked to provide a rationale for why or why not. Annotations involved exploration of the domain and lateral reading—looking at what other reliable sources have written about the domain. We publicly release these annotations alongside annotator rationales. [^14] Following acceptance.
Figure 5: GNN discovery performance vs. classifier confidence reveals that the Partial F1 metric is precision bounded.